A mass m1 = 5.8 kg rests on a frictionless table and connected by a massless string to another mass m2 = 4.5 kg. A force of magnitude F = 30 N pulls m1 to the left a distance d = 0.86 m.

What is the NET work done on m1?
Answer #1

Total weight=5.8+4.5=10.3 kg

Net acceleration= Force/mass=30/10.3=2.912 m/s2

Force on m1=mass*acceleration=5.8*2.912=16.89 N

Work done=Force*distance moved=16.89*0.86

                                                                =14.53 J
Answer #2

Work done is Force * displacement or 30*0.86 J or 25.8 J
Answer #3

30*086=.5*(5.8+4.5)*v^2 solve for v

NET work done on m1= .5*(5.8)*v^2 use v from above